A swimming pool is 7 meters wide and 18 meters long. At $15 per meter,
how much will fencing around the outside of the pool cost?
I tried to find the area of the triangle but that didn't come out with the right answer.

Answer by Alan3354(69443)

It's not the area that's fenced, it's the perimeter. And it's not a triangle.
P = 2(L + W) = 2(7 + 18)
P = 50 meters
50*$15 = $750